How do you change this fraction to an improper fraction?

13 2/5

Answer by stanbon(75887)

13 2/5
Change the 13 to a fraction with 5 as its denominator, as follows:
5(13)/5 = 65/5
Add that to 2/5 to get 67/5
The way this is usually done is [13(5)+2]/5 = 67/5
Cheers,
Stan H.